# MemoryFormer: Minimize Transformer Computation by Removing Fully-Connected Layers

### v1

@@ Abstract
In order to reduce the computational complexity of large language models, great efforts have been made to to improve the efficiency of transformer models such as linear attention and flash-attention. However, the model size and corresponding computational complexity are constantly scaled up in pursuit of higher performance. In this work, we present MemoryFormer, a novel transformer architecture which significantly reduces the computational complexity (FLOPs) from a new perspective. We eliminate nearly all the computations of the transformer model except for the necessary computation required by the multi-head attention operation. This is made possible by utilizing an alternative method for feature transformation to replace the linear projection of fully-connected layers. Specifically, we first construct a group of in-memory lookup tables that store a large amount of discrete vectors to replace the weight matrix used in linear projection. We then use a hash algorithm to retrieve a correlated subset of vectors dynamically based on the input embedding. The retrieved vectors combined together will form the output embedding, which provides an estimation of the result of matrix multiplication operation in a fully-connected layer. Compared to conducting matrix multiplication, retrieving data blocks from memory is a much cheaper operation which requires little computations. We train MemoryFormer from scratch and conduct extensive experiments on various benchmarks to demonstrate the effectiveness of the proposed model.
@@ 1 Introduction
Since the amount of computation produced by the hash operation is negligible, the absolute majority of the computations now results from the matrix-multiplications within the multi-head attention. Figure 1 demonstrates the comparison of the computational complexity of one block between the proposed MemoryFormer (red) and the baseline transformer (blue). The MemoryFormer block only requires $\sim$19of the FLOPs compared with the baseline transformer block when sequence length $s=2048$ and hidden size $d=2048$. The effect of FLOPs-reduction is going to be more significant as the model size is scaled up. Replacing all fully-connected layers with Memory Layers allows us to trade the memory resources for less computational complexity.
@@ 2.1 Background
Let $\mathbf{x}\in\mathbb{R}^{d}$ be a row vector representing any token embedding, a fully-connected layer parameterized by weight matrix $\mathbf{W}\in\mathbb{R}^{d\times h}$ applies a linear projection to $\mathbf{x}$ formulated as $\mathbf{y}=\mathbf{x}\mathbf{W}$, where $\mathbf{y}\in\mathbb{R}^{h}$ is the output token embedding. For a sequence composed of $s$ tokens, this would become a matrix multiplication in Eq. 1 with computational complexity of $\mathcal{O}(sdh)$.
$\mathbf{Y}=\mathbf{X}\mathbf{W},~{}~{}~{}\mathbf{X}\in\mathbb{R}^{s\times d},~{}\mathbf{y}\in\mathbb{R}^{s\times h}.\vspace{-0.4mm}$ (1)
@@
In this paper, we use normal-font lowercase letters to denote scalars (e.g. $d$ and $h$), bold-font lowercase letters to denote vectors (e.g. $\mathbf{x}$ and $\mathbf{y}$), and bold-font uppercase letters to denote matrices (e.g. $\mathbf{W}$). We use the notation $[\cdot]_{i}$ to represent the $i$-th row of a matrix, and also use $[\cdot]_{i}$ to represent the $i$-th entry of a vector.
@@ 2.2 Compute-less Locality-Sensitive Hashing
Firstly, we construct an in-memory hash table parameterized by the matrix $\mathbf{T}\in\mathbb{R}^{2^{d}\times h}$ which stores $2^{d}$ vectors $[\mathbf{T}]_{i}\in\mathbb{R}^{h}$. Traditional LSH functions, such as Hyperplane-LSH [5], incorporate multiple linear projections to generate the hash code for table indexing. To avoid any unnecessary computations, we utilize a much simpler LSH function to generate the hash code $h(\mathbf{x})$. Specifically, the process of hashing and retrieving the result from table $\mathbf{T}$ is formulated as:
@@
$\text{sign}([\mathbf{x}]_{i})=\left\{\begin{array}[]{lr}-1,~{}\text{if}~{}[\mathbf{x}]_{i}<0,\\ ~{}~{}~{}1,~{}\text{if}~{}[\mathbf{x}]_{i}\geq 0,\\ \end{array}\right.$ (5)
@@
Since $\mathbf{z}_{k}$ has a smaller bit width after binarization and thus the corresponding hash table $\mathbf{T}_{k}$ would comsume less memory space for storage. The space complexity of Eq. 9 is $\mathcal{O}(K2^{\tau}h)$. When $d=h=512,\tau=8,K=64$ and data type is float16, the storage required by all $K$ hash tables is $\sim 16$ MegaBytes(MB). Figure 2 is a simple demonstration of the proposed locality-sensitive hashing.
@@ 2.3 Memory Layer
So far, the above-mentioned formulation is able to simulate the forward pass of fully-connected layer. And the values store in the hash tables can be updated via back-propagation. The derivative of the loss function $L$ with respect to the hash table is $\frac{\partial L}{\partial[\mathbf{T}_{k}]_{h(\mathbf{z}_{k})}}=\frac{\partial L}{\partial\mathbf{\hat{y}}}\frac{\partial\mathbf{\hat{y}}}{\partial[\mathbf{T}_{k}]_{h(\mathbf{z}_{k})}}$. However, the input vector $\mathbf{x}$ is unable to have gradient since it’s hashed to multiple integers $h(\mathbf{z}_{k})$ used as the index number for retrieval, which is a non-differentiable operation. If we can reformulate Eq. 9 as $\hat{\mathbf{y}}=\sum_{k=1}^{K}~{}p(\mathbf{z}_{k})\cdot[\mathbf{T}_{k}]_{h(\mathbf{z}_{k})}$ to add a coefficient $p(\mathbf{z}_{k})$ to weight each retrieved item, where $p(\mathbf{z}_{k})$ is a function of the variable $\mathbf{z}_{k}$, the gradients can be back-propagated to the input $\mathbf{x}$ via $[\mathbf{T}_{k}]_{h(\mathbf{z}_{k})}$.
As we can observe from Figure 2, many sub-vectors with various directions and amplitudes can still be hashed to the same bucket as long as their signs are identical, but the angle between the bucket’s representative binary vector (each entry is either 1 or -1) and these sub-vectors are different, which is defined by the cosine value cos($\mathbf{z}_{k},\text{sign}(\mathbf{z}_{k})$). We use a scaled cosine similarity, which takes into account both the direction and amplitude of $\mathbf{z}_{k}$, to measure the relevance between $\mathbf{z}_{k}$ and its corresponding hash bucket $h(\mathbf{z}_{k})$:
@@
where $\langle$,$\rangle$ , ⟩ computes the inner-product of two vectors. Considering all the $2^{\tau}$ buckets that $\mathbf{z}_{k}$ is possibly hashed to in the lookup table $\mathbf{T}_{k}$ simultaneously, we define the probability that $\mathbf{z}_{k}$ is specifically mapped to the $h(\mathbf{z}_{k})$-th hash bucket:
@@
The left part of Figure 3 illustrates the schematic of the Memory Layer. For a sequence composed of $s~{}d$-dimensional tokens, and the dimensionality of output embeddings is $h$, the computation complexity of a Memory Layer in Eq. 13 is $\mathcal{O}(s(\tau+h)K)\approx\mathcal{O}(\frac{sdh}{\tau})$. This is one order of magnitude smaller than the fully-connected layer which is $\mathcal{O}(sdh)$ when $\tau=10$.
According to Eq. 13, we can compute the derivative of the loss function $L$ with respect to both the hash tables and the input vector as follows:
@@ 2.4 Architecture of MemoryFormer
We follow the generic design paradigm of the standard transformer architecture [25] using N stacked blocks to build the MemoryFormer. The right part of Figure 3 depicts one building block.
@@
In MemoryFormer we use the Memory Block to replace the Feed-Forward Network used in standard transformers. The Memory Block is composed of 2 consecutive Memory Layers, each of which is preceded by a Norm($\cdot$) layer. The norm layer is vital by setting a zero-mean distribution for the input embedding before the hashing operation, and thus the sign function in Eq. 5 can generate -1 and +1 evenly. Therefore, the output of Eq. 6 can have an uniform distribution so that every bucket in the hash table will be retrieved with equiprobability.
@@
[…]
@@
[…]
@@ 3.1 Evaluation Across Different Scales
We choose Pythia-70M, Pythia-160M, Pythia-410M as the baseline models upon which we build our MemoryFormers. Specifically, MemoryFormer-tiny has the same hidden size and number of layers as Phythia-70M, MemoryFormer-small has the same hidden size and number of layers as Phythia-160M and MemoryFormer-base has the same hidden size and number of layers as Phythia-410M. As for the hyper-parameter of Memory Layer, we fix the value of $\tau$ to be 8, while the number of hash tables $K$ is 64, 96 and 128 respectively for MemoryFormer-tiny, -small and -base model. Notably, considering the sparsity of gradients of the hash tables, we set the learning rate to be 3 times of the baseline learning rate used by the corresponding Pythia model. This is ablated in Sec. 3.3. It’s worth noting that, the only one fully-connected layer in the MemoryFormer is the classifier head.
Tab. 1 reports both the computational complexity and the evaluation results of our models compared to baseline. The FLOPs is calculated for one transformer block with the input sequence length of 2048. The experiment results show that MemoryFormer has the minimum computation except for the necessary computation of self-attention. Across all three different model sizes, we achieve better average accuracy on the benchmark than the Pythia baseline. This suggests that the proposed method is able to greatly reduce the computational complexity without compromising the performance.
@@ 3.2 Comparison with Efficient Transformers
We also compare our models with existing efficient transformer methods to show the superiority of MemoryFormer in both performance and efficiency. We choose Pythia-410M as the baseline, and replace the multi-head attention module of Pythia with the ones proposed by Linformer [26], Cosformer [22], and Performer [8], respectively. Tab. 2 demonstratse the experiment results on the benchmark and the inference FLOPs of each model. We measure the FLOPs using one transformer block with the sequence length of 2048. As shown in Tab. 2, these efficient attention methods can obtain FLOPs-reduction but with considerable performance degradation, while MemoryFormer  significantly eliminates the computations and gains better performance. On the other hand, even though existing efficient transformer methods can reduce the computation cost of the self-attention operation, yet we can observe from Tab. 2 that the majority of the computation originates from the fully-connected layers in both the MHA and FFN module as discussed previously. The linear projection operation accounts for the biggest part of the workload when the sequence length is not extremely large in most of the practical scenarios. Utilizing Memory Layer in the embedding space to replace FC layers does provide a new solution to minimize the FLOPs of LLMs.
@@ 3.3 Ablation Study
In Eq. 8 we split an input embedding $\mathbf{x}$ in to K sub-vectors to avoid the explosive growth of memory usage of the hash tables. We study the model performance with different ($\tau,K$) combination by controlling the model hidden size to be the same. We report the experiment results in Tab. 4 along with the computation in FLOPs of the corresponding Memory Layer with sequence length of 2048. As is shown in Tab. 4, as the bit width of $\mathbf{z}$ increases, the model performance continues to grow due to the exponentially enlarged capacity of hash tables. However, the memory consumption of the Memory Layer also increases drastically and its computational complexity soon reaches the lower bound. To trade off between the efficiency and memory usage, we conjecture that $\tau$ = $8$ is a good option for MemoryFormer.
@@
Table 5: Different expanding bits of Memory Block. #Expanding Bit$=\tau^{\prime}-\tau$ denotes the number of extra bit of $\mathbf{z}_{k}$ after expansion. Memory Size denotes the storage space required by Memory Block.
@@
From Eq. 14 we can observe that the gradients of the hash table are sparse during backward propagation. Some buckets in the hash table might not get updated in one training step. We conjecture that larger learning rate will help remedy the lack-of-gradients situation. We train a MemoryFormer-tiny for 8000 steps with different LR and report the PPL of validation set in Tab. 4. We use initial LR=1e-3 as the baseline learning rate following the settings of Pythia-70M. The best performance is achieved with the learning rate 3 times of the baseline.
@@
As mentioned in Sec. 2.4, in order to increase the model capacity, we enlarge the dimensionality of the output embedding of the first layer of the Memory Block, which consequently expands the bit-width of the the sub-vector $\mathbf{z}_{k}$ in the second layer of the Memory Block. We use MemoryFormer-tiny with hidden size $d=512$, bit-width $\tau=8$ and number of hash tables $K=64$ as the baseline model, and report the perplexity results of using different number of expanding bits after 8000 training steps in Tab. 5. We use $\mathbf{T}_{k}^{M1}$ to denote the hash tables of the 1st layer of the Memory Block, $\mathbf{T}_{k}^{M2}$ to denote the hash tables of the 2nd layer, and use $\tau^{\prime}$ to denote the bit-width of $\mathbf{z}_{k}$ in the second layer after expansion. As shown in Tab. 5, the validation PPL continuously decreases as $\tau^{\prime}$ gets larger. However, the memory space consumed by the hash tables keeps increasing exponentially. Therefore, we choose 2 as the number of expanding bit in the Memory Block for the trade-off between the space complexity and performance.
@@
We also mentioned in Sec. 2.4 that all activation functions are discarded in the MemoryFormer. We trained a MemoryFormer-tiny on PILE dataset, where we insert a GeLU layer between the two consecutive Memory Layers of the Memory Block for each building block, and report the testing scores on multiple tasks. As shown in Tab. 6, adding an extra GeLU into the Memory Block leads to almost identical result to the baseline MemoryFormer.

### v2

@@ Abstract
In order to reduce the computational complexity of large language models, great efforts have been made to to improve the efficiency of transformer models such as linear attention and flash-attention. However, the model size and corresponding computational complexity are constantly scaled up in pursuit of higher performance. In this work, we present MemoryFormer, a novel transformer architecture which significantly reduces the computational complexity (FLOPs) from a new perspective. We eliminate nearly all the computations of the transformer model except for the necessary computation required by the multi-head attention operation. This is made possible by utilizing an alternative method for feature transformation to replace the linear projection of fully-connected layers. Specifically, we first construct a group of in-memory lookup tables that store a large amount of discrete vectors to replace the weight matrix used in linear projection. We then use a hash algorithm to retrieve a correlated subset of vectors dynamically based on the input embedding. The retrieved vectors combined together will form the output embedding, which provides an estimation of the result of matrix multiplication operation in a fully-connected layer. Compared to conducting matrix multiplication, retrieving data blocks from memory is a much cheaper operation which requires little computations. We train MemoryFormer from scratch and conduct extensive experiments on various benchmarks to demonstrate the effectiveness of the proposed model. Code is available at https://github.com/ningding-o/MemoryFormer.
@@ 1 Introduction
Since the amount of computation produced by the hash operation is negligible, the absolute majority of the computations now results from the matrix-multiplications within the multi-head attention. Figure˜1 demonstrates the comparison of the computational complexity of one block between the proposed MemoryFormer (red) and the baseline transformer (blue). The MemoryFormer block only requires $\sim$19% of the FLOPs compared with the baseline transformer block when sequence length $s=2048$ and hidden size $d=2048$. The effect of FLOPs-reduction is going to be more significant as the model size is scaled up. Replacing all fully-connected layers with Memory Layers allows us to trade the memory resources for less computational complexity.
@@ 2.1 Background
Let $\mathbf{x}\in\mathbb{R}^{d}$ be a row vector representing any token embedding, a fully-connected layer parameterized by weight matrix $\mathbf{W}\in\mathbb{R}^{d\times h}$ applies a linear projection to $\mathbf{x}$ formulated as $\mathbf{y}=\mathbf{x}\mathbf{W}$, where $\mathbf{y}\in\mathbb{R}^{h}$ is the output token embedding. For a sequence composed of $s$ tokens, this would become a matrix multiplication in Eq.˜1 with computational complexity of $\mathcal{O}(sdh)$.
$\mathbf{Y}=\mathbf{X}\mathbf{W},~~~\mathbf{X}\in\mathbb{R}^{s\times d},~\mathbf{y}\in\mathbb{R}^{s\times h}.\vskip-1.13809pt$ (1)
@@
In this paper, we use normal-font lowercase letters to denote scalars (e.g. $d$ and $h$), bold-font lowercase letters to denote vectors (e.g. $\mathbf{x}$ and $\mathbf{y}$), and bold-font uppercase letters to denote matrices (e.g. $\mathbf{W}$). We use the notation $\cdot]_{i}$ to represent the $i$-th row of a matrix, and also use $\cdot]_{i}$ to represent the $i$-th entry of a vector.
@@ 2.2 Compute-less Locality-Sensitive Hashing
Firstly, we construct an in-memory hash table parameterized by the matrix $\mathbf{T}\in\mathbb{R}^{2^{d}\times h}$ which stores $2^{d}$ vectors $\mathbf{T}]_{i}\in\mathbb{R}^{h}$. Traditional LSH functions, such as Hyperplane-LSH [5], incorporate multiple linear projections to generate the hash code for table indexing. To avoid any unnecessary computations, we utilize a much simpler LSH function to generate the hash code $h(\mathbf{x})$. Specifically, the process of hashing and retrieving the result from table $\mathbf{T}$ is formulated as:
@@
[…]
@@
Since $\mathbf{z}_{k}$ has a smaller bit width after binarization and thus the corresponding hash table $\mathbf{T}_{k}$ would comsume less memory space for storage. The space complexity of Eq.˜9 is $\mathcal{O}(K2^{\tau}h)$. When $d=h=512,\tau=8,K=64$ and data type is float16, the storage required by all $K$ hash tables is $\sim 16$ MegaBytes(MB). Figure˜2 is a simple demonstration of the proposed locality-sensitive hashing.
@@ 2.3 Memory Layer
So far, the above-mentioned formulation is able to simulate the forward pass of fully-connected layer. And the values store in the hash tables can be updated via back-propagation. The derivative of the loss function $L$ with respect to the hash table is $\frac{\partial L}{\partial[\mathbf{T}_{k}]_{h(\mathbf{z}_{k})}}=\frac{\partial L}{\partial\mathbf{\hat{y}}}\frac{\partial\mathbf{\hat{y}}}{\partial[\mathbf{T}_{k}]_{h(\mathbf{z}_{k})}}$. However, the input vector $\mathbf{x}$ is unable to have gradient since it’s hashed to multiple integers $h(\mathbf{z}_{k})$ used as the index number for retrieval, which is a non-differentiable operation. If we can reformulate Eq.˜9 as $\hat{\mathbf{y}}=\sum_{k=1}^{K}~p(\mathbf{z}_{k})\cdot[\mathbf{T}_{k}]_{h(\mathbf{z}_{k})}$ to add a coefficient $p(\mathbf{z}_{k})$ to weight each retrieved item, where $p(\mathbf{z}_{k})$ is a function of the variable $\mathbf{z}_{k}$, the gradients can be back-propagated to the input $\mathbf{x}$ via $\mathbf{T}_{k}]_{h(\mathbf{z}_{k})}$.
[…]
@@
where $\langle,\rangle$ computes the inner-product of two vectors. Considering all the $2^{\tau}$ buckets that $\mathbf{z}_{k}$ is possibly hashed to in the lookup table $\mathbf{T}_{k}$ simultaneously, we define the probability that $\mathbf{z}_{k}$ is specifically mapped to the $h(\mathbf{z}_{k})$-th hash bucket:
@@
The left part of Figure˜3 illustrates the schematic of the Memory Layer. For a sequence composed of sds~d-dimensional tokens, and the dimensionality of output embeddings is $h$, the computation complexity of a Memory Layer in Eq.˜13 is $\mathcal{O}(s(\tau+h)K)\approx\mathcal{O}(\frac{sdh}{\tau})$. This is one order of magnitude smaller than the fully-connected layer which is $\mathcal{O}(sdh)$ when $\tau=10$.
According to Eq.˜13, we can compute the derivative of the loss function $L$ with respect to both the hash tables and the input vector as follows:
@@ 2.4 Architecture of MemoryFormer
We follow the generic design paradigm of the standard transformer architecture [25] using N stacked blocks to build the MemoryFormer. The right part of Figure˜3 depicts one building block.
@@
In MemoryFormer we use the Memory Block to replace the Feed-Forward Network used in standard transformers. The Memory Block is composed of 2 consecutive Memory Layers, each of which is preceded by a Norm($\cdot$) layer. The norm layer is vital by setting a zero-mean distribution for the input embedding before the hashing operation, and thus the sign function in Eq.˜5 can generate -1 and +1 evenly. Therefore, the output of Eq.˜6 can have an uniform distribution so that every bucket in the hash table will be retrieved with equiprobability.
@@
$\mathbf{Q}$ $=\text{MemoryLayer${}_{Q}$}(\mathbf{X}),~~\mathbf{K}=\text{MemoryLayer${}_{K}$}(\mathbf{X}),~~\mathbf{V}=\text{MemoryLayer${}_{V}$}(\mathbf{X}),$
@@
$\mathbf{Y}$ $=\mathbf{Z}+\text{MemoryLayer${}_{2}$}(\text{Norm}(\text{MemoryLayer${}_{1}$}(\text{Norm}(\mathbf{X})))).$
@@ 3.1 Evaluation Across Different Scales
We choose Pythia-70M, Pythia-160M, Pythia-410M as the baseline models upon which we build our MemoryFormers. Specifically, MemoryFormer-tiny has the same hidden size and number of layers as Phythia-70M, MemoryFormer-small has the same hidden size and number of layers as Phythia-160M and MemoryFormer-base has the same hidden size and number of layers as Phythia-410M. As for the hyper-parameter of Memory Layer, we fix the value of $\tau$ to be 8, while the number of hash tables $K$ is 64, 96 and 128 respectively for MemoryFormer-tiny, -small and -base model. Notably, considering the sparsity of gradients of the hash tables, we set the learning rate to be 3 times of the baseline learning rate used by the corresponding Pythia model. This is ablated in Sec.˜3.3. It’s worth noting that, the only one fully-connected layer in the MemoryFormer is the classifier head.
Tab.˜1 reports both the computational complexity and the evaluation results of our models compared to baseline. The FLOPs is calculated for one transformer block with the input sequence length of 2048. The experiment results show that MemoryFormer has the minimum computation except for the necessary computation of self-attention. Across all three different model sizes, we achieve better average accuracy on the benchmark than the Pythia baseline. This suggests that the proposed method is able to greatly reduce the computational complexity without compromising the performance.
@@ 3.2 Comparison with Efficient Transformers
We also compare our models with existing efficient transformer methods to show the superiority of MemoryFormer in both performance and efficiency. We choose Pythia-410M as the baseline, and replace the multi-head attention module of Pythia with the ones proposed by Linformer [26], Cosformer [22], and Performer [8], respectively. Tab.˜2 demonstratse the experiment results on the benchmark and the inference FLOPs of each model. We measure the FLOPs using one transformer block with the sequence length of 2048. As shown in Tab.˜2, these efficient attention methods can obtain FLOPs-reduction but with considerable performance degradation, while MemoryFormer  significantly eliminates the computations and gains better performance. On the other hand, even though existing efficient transformer methods can reduce the computation cost of the self-attention operation, yet we can observe from Tab.˜2 that the majority of the computation originates from the fully-connected layers in both the MHA and FFN module as discussed previously. The linear projection operation accounts for the biggest part of the workload when the sequence length is not extremely large in most of the practical scenarios. Utilizing Memory Layer in the embedding space to replace FC layers does provide a new solution to minimize the FLOPs of LLMs.
@@ 3.3 Ablation Study
In Eq.˜8 we split an input embedding $\mathbf{x}$ in to K sub-vectors to avoid the explosive growth of memory usage of the hash tables. We study the model performance with different ($\tau,K$) combination by controlling the model hidden size to be the same. We report the experiment results in Tab.˜4 along with the computation in FLOPs of the corresponding Memory Layer with sequence length of 2048. As is shown in Tab.˜4, as the bit width of $\mathbf{z}$ increases, the model performance continues to grow due to the exponentially enlarged capacity of hash tables. However, the memory consumption of the Memory Layer also increases drastically and its computational complexity soon reaches the lower bound. To trade off between the efficiency and memory usage, we conjecture that $\tau$ = $8$ is a good option for MemoryFormer.
@@
[…]
@@
From Eq.˜14 we can observe that the gradients of the hash table are sparse during backward propagation. Some buckets in the hash table might not get updated in one training step. We conjecture that larger learning rate will help remedy the lack-of-gradients situation. We train a MemoryFormer-tiny for 8000 steps with different LR and report the PPL of validation set in Tab.˜4. We use initial LR=1e-3 as the baseline learning rate following the settings of Pythia-70M. The best performance is achieved with the learning rate 3 times of the baseline.
@@
As mentioned in Sec.˜2.4, in order to increase the model capacity, we enlarge the dimensionality of the output embedding of the first layer of the Memory Block, which consequently expands the bit-width of the the sub-vector $\mathbf{z}_{k}$ in the second layer of the Memory Block. We use MemoryFormer-tiny with hidden size $d=512$, bit-width $\tau=8$ and number of hash tables $K=64$ as the baseline model, and report the perplexity results of using different number of expanding bits after 8000 training steps in Tab.˜5. We use $\mathbf{T}_{k}^{M1}$ to denote the hash tables of the 1st layer of the Memory Block, $\mathbf{T}_{k}^{M2}$ to denote the hash tables of the 2nd layer, and use $\tau^{\prime}$ to denote the bit-width of $\mathbf{z}_{k}$ in the second layer after expansion. As shown in Tab.˜5, the validation PPL continuously decreases as $\tau^{\prime}$ gets larger. However, the memory space consumed by the hash tables keeps increasing exponentially. Therefore, we choose 2 as the number of expanding bit in the Memory Block for the trade-off between the space complexity and performance.
@@
We also mentioned in Sec.˜2.4 that all activation functions are discarded in the MemoryFormer. We trained a MemoryFormer-tiny on PILE dataset, where we insert a GeLU layer between the two consecutive Memory Layers of the Memory Block for each building block, and report the testing scores on multiple tasks. As shown in Tab.˜6, adding an extra GeLU into the Memory Block leads to almost identical result to the baseline MemoryFormer.